A car passes a landmark on a highway traveling at a constant rate of 45 kilometers per hour. An hour later, a second car passes the same landmark traveling in the same direction at 60 kilometers per hour. How much time after the second car passes the landmark will it overtake the first car?
Answer by mananth(16946)

the distance traveled by both is same when one catches up with the other.
Let them meet at distance x
Car I speed= 45 mph
Car II speed= 60 mph

Time difference between the the two when they cross landmark = 1.00 hour

Time taken by Car I = X / 45
Time taken by Car II = X / 60

X / 45 - X / 60 = 1

LCD - 180
Multiply equation by 180
180/45x-180/60x=1*180
4 x - 3 x = 180
1 x = 180
/ 1
x= 180 miles
Car II speed= 60 mph
Distance = 180 miles
Time taken = 180 / 60
Time taken = 3 hour

They will be together after 3 hours

m.ananth@hotmail.ca